A uniform magnetic field of magnitude 0.20 T exists in space from east to west. With what speed should a particle of mass 0.0.10 g and having a charge 1.0 × 10^-5 C be projected from south to north so that it moves with a uniform velocity?

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. B = 0.20 T, υ = ? m = 0.010g = 10^–5 kg, q = 1 × 10^–5 C Force due to magnetic field = Gravitational force of attraction So, qυB = mg ⇒ 1 × 10^–5 × υ × 2 × 10^–1 = 1 × 10^–5 × 9.8 ⇒ υ = 9.8 * 10^-5/2 * 10^-6 = 49 m/s.